# Evaluating Text-to-Image Synthesis with a Conditional Fréchet Distance

### v1

@@ 3 Methodology
Table 1: Text-to-image model ranking and scores by statistical models (FID, $\text{FD}_{\text{\tiny{DINOv2}}}$, CLIP score, CMMD, and cFreD) and models that were trained with human preference (Aesthetic Score, ImageReward, HPSv2, and MPS) on HPDv2 test set. Best results in bold, second best underlined
3.1 Conditional Fréchet Distance
$Q_{y|x}$ $=\mathcal{N}(\mu_{y|x},\,\Sigma_{yy|x}),$ (1)
$Q_{\hat{y}|x}$ $=\mathcal{N}(\mu_{\hat{y}|x},\,\Sigma_{\hat{y}\hat{y}|x}),$
$\mathrm{cFreD}$ $=\mathbb{E}_{x}\Bigg{[}\|\mu_{y|x}-\mu_{\hat{y}|x}\|^{2}$
$\mathrm{cFreD}=$ $\|\mu_{y}-\mu_{\hat{y}}\|^{2}+\mathrm{Tr}\Big{[}(\Sigma_{yx}-\Sigma_{\hat{y}x})\Sigma_{xx}^{-1}(\Sigma_{xy}-\Sigma_{x\hat{y}})\Big{]}$
4 Experiment Settings
We evaluate the alignment of our proposed metric $\mathrm{cFreD}$ with human preferences and compare it against standard evaluation metrics. In this section we introduce the experimental setup (Sec. 4.1), and report the results (Sec. 4.2).
@@ 4 Experiment Settings
$\bm{\rho}^{2}$ 0.70 0.70 0.12 0.54 0.73 0.43 0.81 0.83 0.65
Table 2: Text-to-image model ranking and scores by statistical metrics (FID, $\text{FD}_{\text{\tiny{DINOv2}}}$, CLIP score, CMMD, and $\mathrm{cFreD}$) and models that were trained with human preference (Aesthetic Score, ImageReward, and MPS) on Parti-Prompt. Best results in bold, second best underlined
4.1 Experimental Setup
Dataset. To evaluate how correlated our metric is with human preferences, we need rankings of generated images from different models using the same prompt. We use HPDv2 test set [73] and Parti Prompts Arena. HPDv2 [73] test set contains nine different generated images (each from a different model) and one real image from COCO [6] for each prompt. Parti Prompts Arena [49] has 1,600 generated images, each from four different models. We gather the scattered Parti Prompts data on the web and consolidate it into a unified version. However, since the real image does not consistently receive the highest human preference, we took the most preferred image for each prompt and designated it as the reference image. To evaluate recent models, we randomly selected 1,000 prompts from the COCO train and validation sets, ensuring these prompts were not part of the HPDv2 training and testing sets. For each prompt, we generated images using nine models listed on the Arena Leaderboard [1]. The original COCO images serve as reference images in this evaluation.
Metric Comparison & Evaluation Setting. We compare $\mathrm{cFreD}$ against four standard statistical metrics (FID [30], $\text{FD}_{\text{\footnotesize{DINOv2}}}$ [67], CLIPScore [29], and CMMD [31]). We also compare with four models that were trained with human preference data (Aesthetic Score [63], ImageReward [76], HPSv2 [73], and MPS [84]). We evaluate human correlation from both ranking and scoring perspectives. For each metric, we report the scores and rankings of the models and calculate the correlation with human evaluation results. $\mathrm{cFreD}$ uses DINOv2-G/14 111timm/vit_giant_patch14_dinov2.lvd142m for image embedding and the OpenCLIP ConvNext-B Text Encoder 222 convnext_base_w.laion2b_s13b_b82k_augreg for text embeddings.
4.2 Results
Results on HPDv2. Table 1 summarizes the rankings and scores for images generated by various models—including one real image—across multiple evaluation metrics. Notably, $\mathrm{cFreD}$ achieves the highest alignment with human preferences, reaching a correlation of 0.97. Among statistical metrics, $\mathrm{cFreD}$ attains the highest correlation and is comparable to HPSv2 (0.94), a model explicitly trained on human preferences. Given that HPSv2 was trained on the HPSv2 training set, which includes four models from the test set, and employed the same annotators, it inherently encodes specific human preference biases of the same setting. In contrast, $\mathrm{cFreD}$ achieves comparable or superior correlation with human evaluation without any human preference training. These results demonstrate that $\mathrm{cFreD}$ provides more reliable rankings across diverse models compared to standard automatic metrics and metrics trained explicitly on human preference data.
Table 1 also reports the rank accuracy scores on the HPDv2 test set. Among all evaluated metrics, $\mathrm{cFreD}$ achieves the highest rank accuracy (91.1%), highlighting its strong correspondence with human judgments. HPSv2 follows as the second-best metric with an accuracy of 88.9%, while both FID and $\text{FD}_{\text{DINOv2}}$ obtain competitive scores of 86.7%. Overall, although models trained with human preference data tend to align well with human judgments, $\mathrm{cFreD}$ emerges as the most robust and reliable metric.
Result on PartiPrompts Arena. Table 2 presents the rankings and scores of text-to-image models evaluated on the Parti-Prompt Arena using both statistical metrics and human preference-trained models. Among the statistical metrics, $\mathrm{cFreD}$ achieves the highest correlation with human evaluations (0.73), with FID and $\text{FD}_{\text{\tiny{DINOv2}}}$ both reaching a correlation of 0.70. In contrast, the CLIP score shows a very low correlation (0.12) with human judgments. In the human preference trained category, HPSv2 has the strongest alignment, achieving the highest correlation (0.83), followed by ImageReward (0.81) and MPS (0.65). These results highlight that while $\mathrm{cFreD}$ is a robust automatic metric, HPSv2 stands out as the most effective in capturing human evaluation trends in the PartiPrompts Arena.
@@ 4.2 Results
Table 3: Text-to-image model ranking by automatic models(FID, $\text{FD}_{\text{\tiny{DINOv2}}}$, CLIP score, CMMD, and cFreD) and models that were trained with human preference (Aesthetic Score, ImageReward, HPSv2, and MPS) on randomly sampled COCO prompts. Rank Acc. below 0.5 indicates there are more discordant pairs than concordant ones. Best results in bold, second best underlined
@@
Results on COCO.  Table 3 presents an evaluation on the COCO dataset using nine modern text-to-image models, with human preference rankings sourced from the Text-to-Image Leaderboard and expressed as ELO scores. Among statistical metrics (FID, $\text{FD}_{\text{\footnotesize{DINOv2}}}$, CLIP, CMMD, and our proposed $\mathrm{cFreD}$), only $\mathrm{cFreD}$ exhibits a strong correlation with human preferences, achieving a correlation of 0.33 and a non-trivial rank accuracy of 66.67%. This result places $\mathrm{cFreD}$ as the third most aligned metric overall, surpassed only by the human preference–trained metrics ImageReward, HPSv2, and MPS. Notably, all other statistical metrics show considerably weaker alignment with ELO rankings and, as a result, inverted the rankings, resulting in a Rank Acc. below 0.5. These findings highlight that $\mathrm{cFreD}$ is sensitive to both visual fidelity and prompt consistency, reinforcing its value as a practical, training-free alternative for benchmarking text-to-image generation.
4.3 Are we done with FID?
(a) Effect of training data.
(b) Effect of image size.
(c) Effect of model size.
(d) Effect of number of features.
(e) Effect of IN1K Zero-shot accuracy.
4.4 Are we done with InceptionV3?
5 Ablations
To analyze how different image and text models affect $\mathrm{cFreD}$, we experimented across all combinations of 43 different image models and 43 different text models to select the backbones for $\mathrm{cFreD}$. The full list of these models is provided in the appendix. In this section, we specifically provide an in-depth analysis of how different visual encoder attributes impact $\mathrm{cFreD}$’s correlation with human preferences. For each factor examined, we report the average correlation across all possible text encoders. We perform our ablations on the PartiPrompts Arena. Detailed results of these analyses are included in the appendix.
Effect of the size of the pre-training dataset on $\mathrm{cFreD}$. Fig. 4(a) shows the correlation between $\mathrm{cFreD}$ and human preferences as a function of the size of the pre-training dataset for Vision Transformers (ViT). Larger datasets tend to improve alignment with human judgments, with models trained on $\approx$10B images achieving the highest correlation of 0.79. Interestingly, the smallest dataset ($\approx$10M) also demonstrates a strong correlation of 0.78, surpassing both the $\approx$2B (0.76) and $\approx$100M (0.73) categories. This non-monotonic relationship suggests that factors beyond raw data quantity, such as diversity and quality, significantly influence performance. Meanwhile, the $\approx$5B dataset yields the lowest correlation of 0.73. These findings emphasize the importance of both data scale and curation in optimizing text-conditioned image generation metrics.
Effect of image size on $\mathrm{cFreD}$. Fig. 4(b) illustrates the correlation between $\mathrm{cFreD}$ and human preferences across varying input image resolutions. We observe a non-monotonic relationship: increasing resolutions do not consistently yield higher correlations. In particular, an image size of 299$\times$229 (corresponding to the InceptionV3) achieves a high correlation of 0.760, while 518$\times$518 (used in all DINOv2 models) reaches a correlation of 0.743. Smaller resolutions, such as 224$\times$224 and 256$\times$256, still perform competitively at 0.719 and 0.701, respectively. However, beyond 518$\times$518, performance declines (e.g., 0.680 at 896$\times$896 and 0.559 at 1024$\times$1024).
Effect of ViT model size on $\mathrm{cFreD}$. The correlation between $\mathrm{cFreD}$ and human preferences across Vision Transformer (ViT) sizes are presented in Fig. 4(c). Larger models consistently achieve higher correlations, with the Huge category reaching the highest correlation of 0.786, followed closely by Large at 0.785. The smallest model (Small) achieves 0.755, indicating a performance gap compared to larger architectures. Intermediate model sizes (Base ) yield correlations of 0.776, while SO (0.724) and Giant (0.755) show a gradual improvement between these model sizes.
Effect of ViT Feature dimensionality on $\mathrm{cFreD}$.  Fig. 4(d) shows the correlation between $\mathrm{cFreD}$ and human preferences across ViT feature dimensions from 256 to 1664. The relationship is non-monotonic: correlation peaks at 1024 features (0.79), while certain lower-dimensional representations (e.g., 640, 512, and 1280 features at 0.78) also perform strongly. Both very low (256 at 0.67) and very high (1536 at 0.75) feature sizes yield reduced correlations. These findings indicate that while increasing feature dimensionality generally improves alignment with human preferences, an optimal range exists where performance is maximized, suggesting a trade-off between model complexity and effective representation learning.
Effect of Zero-Shot ImageNet Accuracy on $\mathrm{cFreD}$.  Fig. 4(e) shows a boxplot of the correlation between $\mathrm{cFreD}$ and human preferences as a function of zero-shot ImageNet accuracy, evaluated exclusively on image-text pretrained models [55, 63, 18, 21, 75]. Higher zero-shot accuracies generally correspond to stronger correlations with human judgments, though variance exists within each accuracy level. Models with the highest accuracy (81.37%) achieve the highest median correlation, while models at the lower end (51.53%) cluster around lower correlation values. These findings suggest that zero-shot ImageNet performance can serve as a useful indicator of downstream alignment with human preferences despite considerable variability among models with similar accuracy levels.
6 Discussion and Conclusion
In this work, we extend and repurpose Conditional Fréchet Inception Distance ($\mathrm{cFID}$), first introduced in [66], for the evaluation of text-to-image generation. Recognizing key shortcomings in current evaluation practices—namely, reliance on outdated image encoders like InceptionV3 and the lack of explicit text-image alignment measures—we analyzed a range of modern vision and text backbone models. Our experiments show that modern transformer-based architectures (e.g., ViT, DINOv2, CLIP) significantly outperform InceptionV3, resulting in metrics that better align with human preferences. We replace the backbone models with a superior model and introduce $\mathrm{cFreD}$.
Unlike traditional metrics such as FID or IS, which solely focus on image quality, $\mathrm{cFreD}$ explicitly measures semantic consistency between generated images and their corresponding text prompts, achieving higher correlation with human judgments. Additionally, our ablation studies reveal that optimal metric performance depends on balancing model complexity, data diversity, and representation quality, rather than simply scaling model size or dataset scale.
We advocate for $\mathrm{cFreD}$ as a robust, user-friendly metric for text-to-image synthesis evaluation. It delivers immediate, reliable assessments without constant updates or extensive human annotation. By open-sourcing our implementation and unified benchmarks, we aim to promote more systematic evaluations and accelerate generative model research.
There are several promising directions to extend the scope of $\mathrm{cFreD}$. First, future research could explore combining $\mathrm{cFreD}$ with human preference–trained models by incorporating their image encoders as backbones within our metric. This hybrid approach leverages $\mathrm{cFreD}$’s robust semantic alignment and human preference data’s fine-grained insights, offering the best of both. Second, while our current formulation is tailored for text-to-image synthesis, $\mathrm{cFreD}$ is inherently flexible and can be adapted to other conditional evaluation tasks. For example, it could be extended to assess image-to-text (captioning), text-to-video, or audio-to-video generation. Furthermore, $\mathrm{cFreD}$ can be generalized to condition on multiple inputs, such as in models like ControlNet [82] that require simultaneous conditioning on a text prompt and an image to generate a new image.
Appendix A Image and Text Backbone Models
In our experiments, we used 46 different image backbone models and 43 different text backbone models. For vision models, we employed those trained with self-supervised learning, including ViT trained on ImageNet-1k [15], ViT trained on ImageNet-21k [68], MAE [27], DINOv2 [50], MoCOv3 [7], and I-JEPA trained on both ImageNet-1k and ImageNet-22k [2]. We also incorporated image-and-text-aligned pretrained models such as CLIP [55], MetaCLIP [75], DFN-CLIP [18], OpenCLIP [9], DataComp-CLIP [21], ConvNeXT-CLIP [63], EVA02 [19], and SigLIP [80]. Additional image models included SAM-ViT [33] and Inception V3 [69]. For text models, we employed autoencoding models such as RoBERTa [45], BERT [13], ALBERT [39], ModernBERT [72], XLM-RoBERTa [12], as well as sequence-to-sequence models including FLAN-T5 [11] and T5 [56] were used. We also used the text encoders from the aforementioned image-and-text-aligned pretrained models. The complete lists of image and text backbone models used in our experiments are presented in Tables B.1 and B.2, respectively.
Appendix B Analysis on Selecting Image and Text Backbone Models
@@ B.1 Spearman Correlation
Fig. B.1 showcases the heatmap on Parti-Prompts. Notably, certain image models such as ViT-B/16 trained on ImageNet-1K and ViT-H/14 trained on ImageNet-21K show consistently high performance with different text models. In contrast, SAM-ViT-H/16 presents more variability with different text models. While most ViT, DINO, and CLIP-based models demonstrate strong correlations across different text models, MAE models show slightly lower correlations.
Fig. B.2 presents the heatmap on HPDv2. Most models show strong correlations, whereas MAE models and SAM have lower correlations. Regardless of their performance levels, all image models show consistent correlation patterns across different text models.
Fig. B.3 provides the heatmap on randomly selected COCO prompts. The correlation values are significantly lower overall, ranging from approximately 0.00 to 0.30, which is much lower than both the Parti-Prompts and HPDv2 datasets. Most models show lower correlations, while only DINOv2 models consistently demonstrate stronger correlations. Compared to other datasets, this heatmap exhibits more variability, such as OpenCLIP showing a mixture of relatively higher and lower correlations depending on the text models.
@@
Figure B.1: Spearman Correlation Heatmap on Parti-Prompts.
Figure B.2: Spearman Correlation Heatmap on HPDv2.
Figure B.3: Spearman Correlation Heatmap on random COCO prompts.
B.2 Effect of Visual Encoders on HPDv2
@@
B.3 Effect of Visual Encoders on COCO prompts
@@
Figure B.4: Ablation study on the HPDv2 dataset comparing the correlation to human preferences under varying factors: (a) the ViT training dataset, (b) input image size, (c) model capacity, (d) the number of features in the last ViT layer, and (e) zero-shot accuracy on ImageNet-1K.
@@
Figure B.5: Ablation study on randomly sampled COCO prompts comparing the correlation to human preferences under varying factors: (a) the ViT training dataset, (b) input image size, (c) model capacity, (d) the number of features in the last ViT layer, and (e) zero-shot accuracy on ImageNet-1K.
@@
Table B.1: List of image backbone models used in our analysis.
@@
Table B.2: List of text backbone models used in our analysis.

### v2

@@ Appendix B Results on Text-to-Image
Table B.1: Text-to-image model ranking and scores by statistical models (FID, $\text{FD}_{\text{\tiny{DINOv2}}}$, CLIP score, CMMD, and cFreD) and models that were trained with human preference (Aesthetic Score, ImageReward, HPSv2, and MPS) on HPDv2 test set. Best results in bold, second best underlined
@@
$\bm{\rho}^{2}$ - 0.70 0.70 0.12 0.54 0.73 0.43 0.81 0.83 0.65
Table B.2: Text-to-image model ranking and scores by statistical metrics (FID, $\text{FD}_{\text{\tiny{DINOv2}}}$, CLIP score, CMMD, and $\mathrm{cFreD}$) and models that were trained with human preference (Aesthetic Score, ImageReward, and MPS) on Parti-Prompt. Best results in bold, second best underlined
@@
Table B.3: Text-to-image model ranking by automatic models(FID, $\text{FD}_{\text{\tiny{DINOv2}}}$, CLIP score, CMMD, and cFreD) and models that were trained with human preference (Aesthetic Score, ImageReward, HPSv2, and MPS) on randomly sampled COCO prompts. Rank Acc. below 0.5 indicates there are more discordant pairs than concordant ones. Best results in bold, second best underlined
Results on HPDv2. Table B.1 summarizes the rankings and scores for images generated by various models—including one real image—across multiple evaluation metrics. Notably, $\mathrm{cFreD}$ achieves the highest alignment with human preferences, reaching a correlation of 0.97. Among statistical metrics, $\mathrm{cFreD}$ attains the highest correlation and is comparable to HPSv2 (0.94), a model explicitly trained on human preferences. Given that HPSv2 was trained on the HPSv2 training set, which includes four models from the test set, and employed the same annotators, it inherently encodes specific human preference biases of the same setting. In contrast, $\mathrm{cFreD}$ achieves comparable or superior correlation with human evaluation without any human preference training. These results demonstrate that $\mathrm{cFreD}$ provides more reliable rankings across diverse models compared to standard automatic metrics and metrics trained explicitly on human preference data.
Table B.1 also reports the rank accuracy scores on the HPDv2 test set. Among all evaluated metrics, $\mathrm{cFreD}$ achieves the highest rank accuracy (91.1%), highlighting its strong correspondence with human judgments. HPSv2 follows as the second-best metric with an accuracy of 88.9%, while both FID and $\text{FD}_{\text{DINOv2}}$ obtain competitive scores of 86.7%. Overall, although models trained with human preference data tend to align well with human judgments, $\mathrm{cFreD}$ emerges as the most robust and reliable metric.
Result on PartiPrompts Arena. Table B.2 presents the rankings and scores of text-to-image models evaluated on the Parti-Prompt Arena using both statistical metrics and human preference-trained models. Among the statistical metrics, $\mathrm{cFreD}$ achieves the highest correlation with human evaluations (0.73), with FID and $\text{FD}_{\text{\tiny{DINOv2}}}$ both reaching a correlation of 0.70. In contrast, the CLIP score shows a very low correlation (0.12) with human judgments. In the human preference trained category, HPSv2 has the strongest alignment, achieving the highest correlation (0.83), followed by ImageReward (0.81) and MPS (0.65). These results highlight that while $\mathrm{cFreD}$ is a robust automatic metric, HPSv2 stands out as the most effective in capturing human evaluation trends in the PartiPrompts Arena.
@@
Results on COCO.  Table B.3 presents an evaluation on the COCO dataset using nine modern text-to-image models, with human preference rankings sourced from the Text-to-Image Leaderboard and expressed as ELO scores. Among statistical metrics (FID, $\text{FD}_{\text{\footnotesize{DINOv2}}}$, CLIP, CMMD, and our proposed $\mathrm{cFreD}$), only $\mathrm{cFreD}$ exhibits a strong correlation with human preferences, achieving a correlation of 0.33 and a non-trivial rank accuracy of 66.67%. This result places $\mathrm{cFreD}$ as the third most aligned metric overall, surpassed only by the human preference–trained metrics ImageReward, HPSv2, and MPS. Notably, all other statistical metrics show considerably weaker alignment with ELO rankings and, as a result, inverted the rankings, resulting in a Rank Acc. below 0.5. These findings highlight that $\mathrm{cFreD}$ is sensitive to both visual fidelity and prompt consistency, reinforcing its value as a practical, training-free alternative for benchmarking text-to-image generation.
Appendix C Analysis on Selecting Image and Text Backbone Models
@@ C.1 Spearman Correlation
Fig. C.1 showcases the heatmap on Parti-Prompts. Notably, certain image models such as ViT-B/16 trained on ImageNet-1K and ViT-H/14 trained on ImageNet-21K show consistently high performance with different text models. In contrast, SAM-ViT-H/16 presents more variability with different text models. While most ViT, DINO, and CLIP-based models demonstrate strong correlations across different text models, MAE models show slightly lower correlations.
Fig. C.2 presents the heatmap on HPDv2. Most models show strong correlations, whereas MAE models and SAM have lower correlations. Regardless of their performance levels, all image models show consistent correlation patterns across different text models.
Fig. C.3 provides the heatmap on randomly selected COCO prompts. The correlation values are significantly lower overall, ranging from approximately 0.00 to 0.30, which is much lower than both the Parti-Prompts and HPDv2 datasets. Most models show lower correlations, while only DINOv2 models consistently demonstrate stronger correlations. Compared to other datasets, this heatmap exhibits more variability, such as OpenCLIP showing a mixture of relatively higher and lower correlations depending on the text models.
@@
C.2 Effect of Visual Encoders on HPDv2
@@
Figure C.1: Spearman Correlation Heatmap on Parti-Prompts.
Figure C.2: Spearman Correlation Heatmap on HPDv2.
@@
Figure C.3: Spearman Correlation Heatmap on random COCO prompts.
@@
C.3 Effect of Visual Encoders on COCO prompts
@@
Figure C.4: Ablation study on the HPDv2 dataset comparing the correlation to human preferences under varying factors: (a) the ViT training dataset, (b) input image size, (c) model capacity, (d) the number of features in the last ViT layer, and (e) zero-shot accuracy on ImageNet-1K.
@@
Figure C.5: Ablation study on randomly sampled COCO prompts comparing the correlation to human preferences under varying factors: (a) the ViT training dataset, (b) input image size, (c) model capacity, (d) the number of features in the last ViT layer, and (e) zero-shot accuracy on ImageNet-1K.
@@
Table C.1: List of image backbone models used in our analysis.
@@
Table C.2: List of text backbone models used in our analysis.